# Unharmful Backdoor-based Client-side Watermarking in Federated Learning

### v1

@@ 5.3 Effectiveness in Conflicting Scenario
In the scenario with backdoor conflicts, we introduce multiple pairs of watermark conflicts among the benign clients. Specifically, we design the same trigger targeting different labels, thereby creating deliberate conflicts in watermark verification. As presented in Table II, without-Sanitizer is unable to achieve a satisfactory WDR, while the corresponding ASR remains dangerously high. This confirms that conflicts compromise the effectiveness of the watermark by making them being unlearnable, leading to ambiguous contribution claims. Meanwhile, the backdoor attack remains highly effective. After Sanitizer conducts the unharmful transformation for every backdoor independently, the WDR improves to an impressive rate exceeding 98%, which is sufficient for contribution verification. More importantly, the ASR experiences a marked reduction of at least 86% and ACC also remains competitive at a considerable level, with a moderate ACC drop of around 5% induced by Sanitizer. The experimental results across three datasets are consistent with those observed in the non-conflicting setting. Sanitizer reduces the ASR to a near-random guess level while maintaining a high WDR and a robust ACC in the conflicting scenario. What stands out from this experiment is that Sanitizer also effectively transforms the conflicting backdoors into unharmful and conflict-free ones, while mitigating the harmful backdoor effect and preserving the main functionality in scenario with backdoor conflicts.
To distill the findings, Sanitizer consistently demonstrates its general effectiveness in both conflict-free and conflict-present scenarios. Why is it effective? The key lies in its ability to successfully and systematically reverse the embedded trigger pattern in each client-submitted model during the FL process, and eventually confine them to their own unharmful, client-specific subspace. We will provide a more in-depth explanation via visualization in the following subsection.
@@ 5.4 Visual Examples of Trigger Inversion
The previous results demonstrate the overall effectiveness of Sanitizer. In this subsection, we focus on specific clients and employ trigger visualization to further validate the effectiveness of Sanitizer in trigger pattern recovery. All hyperparameters and configurations are maintained at default settings. Table III(b) provides two visual example sets of the original trigger patterns (2nd column) and their eventually recovered versions (3rd column) by Sanitizer from CIFAR10. In the scenario without trigger-output conflicts (Table III(a)), we apply three distinct, commonly used patch-like patterns to the clean data for three participating clients (Client 1, Client 5, and Client 9), setting their desired outputs to different classes (1, 2, and 3) respectively. As we can see, in all cases, the corresponding backdoor trigger is successfully recovered, regardless of whether its real-world purpose is benign or malicious. Specifically, Client 1 achieves a notable trigger recovery and reached a WDR of 100.00% after Sanitizer relearns the reversed trigger-output mapping rule into the unique and unharmful input subspace (4th column). Client 5 and Client 9 demonstrate comparable recovery performance with WDR of 99.96% and 99.28%, respectively.
On the other hand, Table III(b) presents results in the scenario where trigger-output conflicts exist, particularly between Client 1 and Client 5, both of whom use an identical trigger pattern (a white square block) while aiming for different target classes (label 1 and label 2). We observe that Sanitizer is insensitive to the presence of backdoor conflicts, as the trigger patterns recovered by Sanitizer continue to be feasible, and the WDR consistently remains high, with Client 1, Client 5, and Client 9 exhibiting 99.98%, 99.93%, and 98.69%, respectively. This effectiveness is attributed to two facts. First, each client is processed independently before aggregation during the FL process. Second, the final unharmful input subspace is client-specific, effectively resolving the conflict issue. Furthermore, we observe that recovering the trigger from an FL-trained model without Sanitizer appears to be not feasible. Additional results can be found in the Appendix.
@@
Analysis on Effectiveness of Trigger Inversion. First, why can the small extracted backdoor subnet stemming from the client-submitted model be used for trigger inversion? Thanks to the clean unlearning part (Section 4.2), we identify a small set of backdoor-related units within the original client-submitted model, which constitute a backdoor subnet. This backdoor subnet fully encapsulates backdoor knowledge while preserving most of the backdoor functionality, as demonstrated by the low ACC and high WDR of the subnet. For instance, on the CIFAR10 using ResNet18, the backdoor subnet achieves an average ACC of 14.60% and WDR of 99.98%. When applying the trigger recovery and backdoor detection on the extracted backdoor subnet, one can more easily expose the potential backdoor target and enhance the quality of the recovered trigger pattern. Second, why is the round-spread trigger recovery feasible? In each round, the models at the client-side are trained on a sanitized global model from the previous round, and they undergo continuous backdoor attacks for either benign or harmful purposes in this round. The backdoor task exhibits rapid convergence, achieving a high WDR or ASR greater than 99% quickly during local training. Upon receiving each client-submitted model, Sanitizer utilizes the results of reverse engineering from the previous round for that client as the initial values for the reverse optimization in the current round, ultimately facilitating the successful reconstruction of the backdoor trigger for each client. Moreover, a key aspect of Sanitizer’s effectiveness is that the trigger from each client is confined to its own harmless and client-specific input subspace (i.e., being effective only in the unharmful environment), which is directly attributed to the success of the trigger inversion step.
@@ 5.5.1 Efficiency of Sanitizer
As both Sanitizer and the baseline (described in Section 1) aim to achieve the identically expected effectiveness, we primarily evaluate the differences in their efficiency during the FL process, which is a critical concern in real-world FL applications. To this end, we evaluate the efficiency of the two methods on two different architectures, with the other settings kept identical to those in Section 5.3. We use the following primary metrics: (i) the GPU memory utilized GPU_Mem (MB) for processing one client-submitted model during server-side operations and (ii) the average runtime Time (seconds) of server-side operations per round during FL.
Results on ResNet18. Table IV reports the results. What most strikingly emerges from it is that despite both approaches achieve considerable effectiveness, Sanitizer exhibits significantly less overhead in terms of resource consumption and runtime compared to the baseline. Specifically, the baseline consumes 1,322 MB of GPU memory for each client-submitted model, and takes 142.50 seconds for server-side execution per round. On the contrary, Sanitizer drastically reduces the GPU memory usage to 194 MB with an 85% reduction, which suggests that in each round, the server can process more client-submitted models in parallel. Moreover, the runtime optimization achieved by Sanitizer per round over the baseline is also particularly significant, with a 5x speedup in execution time from 142.50 to 28.12 seconds, highlighting the efficiency gains provided by Sanitizer. With the significantly reduced resource and time consumption, Sanitizer still maintains non-trivial effectiveness, closely aligning with the performance of the baseline.
@@
In this subsection, we derive the important insight that Sanitizer demonstrates a significant efficiency advantage while maintaining comparable effectiveness, harmlessness, and fidelity. This is attributed to the fact that Sanitizer performs the round-spread trigger recovery on the extracted small backdoor subnet, two improvements introduced to accelerate the entire FL period. Moreover, the reduction in the GPU memory consumption makes Sanitizer particularly attractive for IP protection in resource-constrained FL environments. Interestingly, this may reflect the scalability of Sanitizer, as already shown in Figure 3 of Section 1, prompting us to further investigate this characteristic, which will be discussed in detail in the subsection 5.5.2.
@@ 5.5.2 Scalability of Sanitizer
To evaluate the scalability of Sanitizer, we examine the total server-side time consumption of two methods during the whole FL process as the number of participating clients increases over a 10-round FL on CIFAR10 using ResNet18. Figure 3 clearly showcases the scalability advantage of Sanitizer when compared to the baseline. Specifically, for the baseline, the time cost increases rapidly, showing near-exponential growth beyond 50 clients. Notably, with 75 clients, the baseline requires over 4,500 seconds to complete, which exceeds the acceptable range we set, and this steep upward trend persists as the number of clients continues to rise. Such scaling behavior suggests the baseline struggles to maintain practical efficiency in large-scale FL scenarios, as the time required to complete sanitization skyrockets when more clients are added. In contrast, Sanitizer exhibits robust scalability across the entire range of clients tested. As shown by the green line, Sanitizer’s time complexity remains consistently low (below 1000 seconds) when scaling from 1 to 150 clients and shows only a moderate linear increase as the number of clients approaches 200. Under the fixed server’s GPU memory, the divergence between the two methods, which initially is fivefold, continues to increase progressively. The stark contrast in time consumption suggests that the improved designs we introduce into Sanitizer allow it to effectively alleviate the severe server-side computation overhead observed in the baseline, making it far more well-suited for real-world deployment in large-scale FL settings, where hundreds or even thousands of clients may be commonplace. Additionally, we believe that the scalability of Sanitizer is also represented in its optimization of GPU resource consumption.
@@
Figure 7 illustrates the comparison of GPU memory consumption between the two methods under the same experimental settings. The baseline exhibits a sharp increase in GPU memory consumption as the number of clients increases. By the time it exceeds 25 clients, the baseline has already surpassed the threshold of 32 GB, a limit for the maximum capacity of the server’s GPU memory in our setting, indicating that a maximum of 25 client-submitted models can be processed concurrently on the server-side. In comparison, Sanitizer is capable of scaling the number of concurrently processed clients beyond 175 while remaining below the threshold, without imposing a significant burden on the server. This is essentially due to the reduced GPU memory consumed by a single backdoor subnet, making Sanitizer well-suited for large-scale FL applications with constrained server-side resources. To sum up, together these results provide a compelling argument for the deployment of Sanitizer in real-world FL environments with a large number of clients.
(a) 5-layer MLP used for FMNIST.
(b) MobileNetV3 used for CIFAR10.
(c) ResNet18 used for TinyImageNet.
@@
5.6 Effectiveness of Sanitizer in Different Neural Architectures
@@
We also observe that Sanitizer remains effective across various architectures. Figure 8 reports the effectiveness of Sanitizer in the scenario with backdoor conflicts, which represents more realistic conditions. Whether applied to traditional fully connected architectures like MLP, classical convolutional architectures such as ResNet18, or lightweight architectures like MobileNetV3, Sanitizer consistently achieves a near-perfect WDR of over 98%, demonstrating its ability to reliably confine the watermarks to a harmless and client-specific input subspace and enable the triggering of counter-intuitive model behavior for contribution demonstration in unharmful, non-conflicting ways. The ASR is reduced from a relatively elevated level to that of random guessing, underscoring that Sanitizer can effectively neutralize malicious backdoors. Additionally, the model’s fidelity is largely preserved, with the ACC typically remaining high. Taken together, this analysis further reveals that Sanitizer is a versatile method that can be effectively applied to different neural architectures in FL.
@@
Figure 9: Efficiency metrics across different settings of round-spread reverse engineering epochs.
Figure 10: Effectiveness metrics across different settings of round-spread reverse engineering epochs.
@@
Figure 11: Efficiency metrics across different settings of the backdoor subnet size (%).
Figure 12: Effectiveness metrics across different settings of the backdoor subnet size (%).
5.7 Evaluation with Different Variants

### v2

@@ 5.3 Effectiveness in Conflicting Scenario
In the scenario with backdoor conflicts, we introduce multiple pairs of watermark conflicts among the benign clients. Specifically, we simulate watermark conflicts by employing identical triggers with different target classes, thereby creating deliberate conflicts in watermark verification. For example, both Clients 1 and 5 use a white square as the trigger; however, Client 1 assigns Class 1 as the target label, while Client 5 assigns Class 2, which will be visualized in Section 5.4. As presented in Table II, without-Sanitizer is unable to achieve a satisfactory WDR, while the corresponding ASR remains dangerously high. This confirms that conflicts compromise the effectiveness of the watermark by making them unlearnable, leading to ambiguous contribution claims. Meanwhile, the backdoor attack remains highly effective. After Sanitizer conducts the harmless transformation for every backdoor independently, the WDR improves to an impressive rate exceeding 96%, which is sufficient for contribution verification. More importantly, the ASR experiences a marked reduction of at least 86%, and ACC also remains competitive at a considerable level, with a moderate ACC drop of around 5% induced by Sanitizer. The experimental results across four datasets are consistent with those observed in the non-conflicting setting. Sanitizer reduces the ASR to a near-random guess level while maintaining a high WDR and a robust ACC in the conflicting scenario. What stands out from this experiment is that Sanitizer also effectively transforms the conflicting backdoors into harmless and conflict-free ones while mitigating the harmful backdoor effect and preserving the main functionality in the scenario with watermark conflicts.
To distill the findings, Sanitizer consistently demonstrates its general effectiveness in both conflict-free and conflict-present scenarios. Why is it effective? The key lies in its ability to successfully and systematically reverse the embedded trigger pattern in each client-submitted model during the FL process and eventually confine them to their own harmless, client-specific input subspace. We will provide a more in-depth explanation via visualization in the following subsection.
@@ 5.4 Visual Examples of Trigger Inversion
The previous results demonstrate the overall effectiveness of Sanitizer. In this subsection, we focus on specific clients and employ trigger visualization to further validate the effectiveness of Sanitizer in trigger pattern recovery. All hyperparameters and configurations are maintained in default settings. Table III(b) provides two visual example sets of the original trigger patterns (2nd column) and their eventually recovered versions (3rd column) by Sanitizer from CIFAR10. In the scenario without trigger-output conflicts (Table III(a)), we apply three distinct, commonly used patch-like patterns to the clean data for three participating clients (Client 1, Client 5, and Client 9), setting their desired outputs to different classes (1, 2, and 3), respectively. As we can see, in all cases, the corresponding backdoor trigger is successfully recovered, regardless of whether its real-world purpose is benign or malicious. Specifically, Client 1 achieves a notable trigger recovery and reaches a WDR of 100.00% after Sanitizer relearns the reversed trigger-output mapping rule into the unique and harmless input subspace (4th column). Client 5 and Client 9 demonstrate comparable recovery performance with WDR of 99.73% and 99.12%, respectively.
@@
On the other hand, Table III(b) presents results in the scenario where trigger-output conflicts exist, particularly between Client 1 and Client 5, both of whom use an identical trigger pattern (a 4$\times$4 white square in the bottom-right corner) while aiming for different target classes (label 1 and label 2). We observe that Sanitizer is insensitive to the presence of backdoor conflicts, as the trigger patterns recovered by Sanitizer continue to be feasible, and the WDR consistently remains high, with Client 1, Client 5, and Client 9 exhibiting 99.98%, 99.93%, and 98.69%, respectively. This effectiveness is attributed to two facts. First, each client is processed independently before aggregation during the FL process. Second, the final harmless input subspace is client-specific, effectively resolving the conflict issue. Furthermore, we observe that recovering the trigger from an FL-trained model without Sanitizer appears to be not feasible. Additional results can be found in the Appendix A.6. Besides, we also report the success rate when only the harmless background is provided (i.e., no trigger is added) during the verification stage. The result is 0%, outputting a randomly fixed label different from the target class. This shows that the trigger’s behavior relies on both the subspace (harmless background) and the trigger content, as the model learns their binding relation during the harmless relearning stage. Using either alone cannot achieve the intended effect in our context.
Analysis on Effectiveness of Trigger Inversion. First, why can the small extracted backdoor subnet stemming from the client-submitted model be used for trigger inversion? Thanks to the clean unlearning part (Section 4.2), we identify a small set of backdoor-related units within the original client-submitted model, which constitute a backdoor subnet. This backdoor subnet fully encapsulates backdoor knowledge while preserving most of the backdoor functionality, as demonstrated by the low ACC and high WDR of the subnet. For instance, on the CIFAR10 using ResNet18, the backdoor subnet achieves an average ACC of 14.60% and WDR of 99.98%. When applying the trigger recovery and backdoor detection on the extracted backdoor subnet, one can more easily expose the potential backdoor target and enhance the quality of the recovered trigger pattern. Second, why is the round-spread trigger recovery feasible? In each round, the models at the client-side are trained on a sanitized global model from the previous round, and they undergo continuous backdoor attacks for either benign or harmful purposes in this round. The backdoor task exhibits rapid convergence, achieving a high WDR or ASR greater than 99% quickly during local training. Upon receiving each client-submitted model, Sanitizer utilizes the results of reverse engineering from the previous round for that client as the initial values for the reverse optimization in the current round, ultimately facilitating the successful reconstruction of the backdoor trigger for each client. Moreover, a key aspect of Sanitizer’s effectiveness is that the trigger from each client is confined to its own harmless and client-specific input subspace (i.e., being effective only in the harmless environment), which is directly attributed to the success of the trigger inversion step.
@@ 5.5.1 Efficiency of Sanitizer
As both Sanitizer and the baseline (described in Section 1) aim to achieve the identically expected effectiveness, we primarily evaluate the differences in their efficiency during the FL process, which is a critical concern in real-world FL applications. To this end, we evaluate the efficiency of the two methods on two different architectures, with the other settings kept identical to those in Section 5.3. We use the following primary metrics: (i) the GPU memory utilized GPU_Mem (MB) for processing one client-submitted model during server-side operations, and (ii) the average runtime Time (s) of server-side operations per round during FL.
Results on ResNet18. Table IV reports the results. What most strikingly emerges from it is that despite both approaches achieving considerable effectiveness, Sanitizer exhibits significantly less overhead in terms of resource consumption and runtime compared to the baseline. Specifically, the baseline consumes 1,322 MB of GPU memory for each client-submitted model and takes 142.50 seconds for server-side execution per round. On the contrary, Sanitizer drastically reduces the GPU memory usage to 194 MB with an 85% reduction, which suggests that in each round, the server can process more client-submitted models in parallel. Moreover, the runtime optimization achieved by Sanitizer per round over the baseline is also particularly significant, with a 5x speedup in execution time from 142.50 to 28.12 seconds, highlighting the efficiency gains provided by Sanitizer. With the significantly reduced resource and time consumption, Sanitizer still maintains non-trivial effectiveness, closely aligning with the performance of the baseline.
@@
In this subsection, we derive the important insight that Sanitizer demonstrates a significant efficiency advantage while maintaining comparable effectiveness, harmlessness, and fidelity. This is attributed to the fact that Sanitizer performs the round-spread trigger recovery on the extracted small backdoor subnet, two improvements introduced to accelerate the entire FL period. Moreover, the reduction in the GPU memory consumption makes Sanitizer particularly attractive for IP protection in resource-constrained FL environments. Interestingly, this may reflect the scalability of Sanitizer, as already shown in Figure 3 of Section 1, prompting us to further investigate this characteristic, which will be discussed in detail in subsection 5.5.2.
@@ 5.5.2 Scalability of Sanitizer
To evaluate the scalability of Sanitizer, we examine the total server-side time consumption of two methods during the whole FL process as the number of participating clients increases over a 10-round FL on CIFAR10 using ResNet18. Figure 3 clearly showcases the scalability advantage of Sanitizer when compared to the baseline. Specifically, for the baseline, the time cost increases rapidly, showing near-exponential growth beyond 50 clients. Notably, with 75 clients, the baseline requires over 4,500 seconds to complete, which exceeds the acceptable range we set, and this steep upward trend persists as the number of clients continues to rise. Such scaling behavior suggests the baseline struggles to maintain practical efficiency in large-scale FL scenarios, as the time required to complete sanitization skyrockets when more clients are added. In contrast, Sanitizer exhibits robust scalability across the entire range of clients tested. As shown by the green line, Sanitizer’s time complexity remains consistently low (below 1,000 seconds) when scaling from 1 to 150 clients and shows only a moderate linear increase as the number of clients approaches 200. Under the fixed server’s GPU memory, the divergence between the two methods, which initially is fivefold, continues to increase progressively. The stark contrast in time consumption suggests that the improved designs we introduce into Sanitizer allow it to effectively alleviate the severe server-side computation overhead observed in the baseline, making it far more well-suited for real-world deployment in large-scale FL settings, where hundreds or even thousands of clients may be commonplace. Additionally, we believe that the scalability of Sanitizer is also represented in its optimization of GPU resource consumption.
Table V: Ablation studies showing the impact of Sanitizer’s two key individual components on efficiency for understanding the contribution of different proposed designs.
GPU_Mem Time
Baseline 1322MB 142.50s
(a) only Subnet Extraction 196MB 113.44s
(b) only Round Spreading 1318MB 69.61s
(a+b) Sanitizer 194MB 28.12s
@@
Figure 7 illustrates the comparison of GPU memory consumption between the two methods under the same experimental settings. The baseline exhibits a sharp increase in GPU memory consumption as the number of clients increases. By the time it exceeds 25 clients, the baseline has already surpassed the threshold of 32 GB, a limit for the maximum capacity of the server’s GPU memory in our setting, indicating that a maximum of 25 client-submitted models can be processed concurrently on the server-side. In comparison, Sanitizer is capable of scaling the number of concurrently processed clients beyond 175 while remaining below the threshold, without imposing a significant burden on the server. This is essentially due to the reduced GPU memory consumed by a single backdoor subnet, making Sanitizer well-suited for large-scale FL applications with constrained server-side resources. To sum up, these results provide a compelling argument for the deployment of Sanitizer in real-world FL environments with a large number of clients.
@@
5.5.3 Ablation Studies
To verify the individual functionality of each key component in Sanitizer and the importance of utilizing them for improving the efficiency, we conduct ablation studies by removing the Subnet Extraction component and Round-spread Optimization component in Sanitizer at a time in default experimental settings. As illustrated in Table V, compared to the baseline, the variant with Subnet Extraction component only (a) cuts the occupied memory per model by approximately 85%, while achieving about a 20% improvement in per-round runtime over the baseline. If we apply Round-spread optimization without extracting subnets (b), it achieves a drastic reduction in runtime by over 51% while maintaining similar memory consumption at 1,318 MB. The final row (a+b) demonstrates the combined effect of both optimizations enabled in our complete Sanitizer, achieving superior performance in both memory usage and runtime compared to either component alone or the baseline. We can derive the following conclusions: (i) Subnet Extraction component is the decisive contributor to GPU memory efficiency by confining the backdoor analysis to a small subnet. Round-spread component alone does not alleviate memory load; (ii) in terms of reducing time overhead, both components demonstrate effectiveness, with the Round-spread component performing better by distributing the inversion optimization across multiple rounds, chiefly contributing to runtime efficiency; (iii) the results demonstrate that employing both components yields the best of both worlds while providing complementary benefits, further indicating their necessity for the success of Sanitizer.
Table VI: Effectiveness comparison of Sanitizer following the Dirichlet Distribution across different concentration parameters $\alpha$, showing Sanitizer’s robust performance across different degrees of data heterogeneity.
Concentration WDR $\uparrow$ ASR $\downarrow$ ACC $\uparrow$
$\alpha=0.7$ 98.61% 17.90% 77.21%
$\alpha=0.8$ 99.22% 14.23% 83.28%
$\alpha=0.9$ 97.13% 12.50% 84.63%
(IID) Sanitizer 99.95% 12.22% 87.18%
5.6 Generalizability of Sanitizer on Different Neural Architectures
@@
We also observe that Sanitizer remains effective across various neural architectures. Figure 8 reports the effectiveness of Sanitizer in the scenario with backdoor conflicts, which represents more realistic conditions. Whether applied to traditional fully connected architectures like MLP, classical convolutional architectures such as ResNet18, or lightweight architectures like MobileNetV3, Sanitizer consistently achieves a near-perfect WDR of over 98%, demonstrating its ability to reliably confine the watermarks to a harmless and client-specific input subspace and enable the triggering of counter-intuitive model behavior for contribution demonstration in harmless, non-conflicting ways. The ASR is reduced from a relatively elevated level to that of random guessing, underscoring that Sanitizer can effectively neutralize malicious backdoors. Additionally, the model’s fidelity is largely preserved, with the ACC typically remaining high.
We enrich our evaluation with ViT [68], even though Transformer-based models are less common in FL due to their computational burden on client devices. Specifically, the core component of ViT, the multi-head self-attention layer, divides the parameter space into multiple heads, where each head specializes in capturing a specific aspect of knowledge. Leveraging this, Sanitizer can identify and extract backdoor-related heads, the associated head channels, and the corresponding hidden dimensions in the MLP module. As shown in Figure 8, Sanitizer works similarly well with a 93.62% success rate for harmless watermark verification and an 11.30% success rate for malicious attacks. Taken together, this analysis further reveals that Sanitizer is a versatile and architecture-agnostic method that can be effectively applied to different neural architectures in FL.
@@
Figure 9: Efficiency and effectiveness metrics across different settings of the round-spread reverse engineering epochs.
@@
Figure 10: Efficiency and effectiveness metrics across different settings of the backdoor subnet size (%).
5.7 Evaluation of Sanitizer under Non-IID Setting
In a more realistic setting for FL, client data distributions frequently exhibit non-independent and identically distributed (non-IID) characteristics, reflecting the heterogeneity of local data among participating clients. To study whether data heterogeneity influences the performance of Sanitizer, we conduct additional experiments and adopt the commonly used Dirichlet distribution [80] to create heterogeneous data partitions across clients with varying degrees of non-IIDness controlled by the concentration parameter $\alpha$. A lower $\alpha$ value indicates higher data skew among clients, and vice versa. We maintain consistent experimental settings and evaluation metrics as detailed in Section 5.2. As shown in Table VI, Sanitizer maintains robust watermark detection performance across different non-IID conditions, with WDR remaining above 97%. It is also observed that as the value of $\alpha$ decreases (i.e., data heterogeneity grows), the ASR climbs. This suggests that skewed data distributions may introduce vulnerabilities that slightly compromise the Sanitizer’s ability to fully neutralize harmful backdoors, enhancing their effectiveness. Additionally, the increased heterogeneity slightly weakens model utility, a result consistent with previous research findings [81, 21], but overall Sanitizer remains resilient in non-IID settings, prioritizing the reliability of the harmless watermarking functionality.
5.8 Comparison with Additional Approaches
To benchmark Sanitizer against more approaches, we next compare it with FedIPR [11], a state-of-the-art client-side FL watermarking method for ownership verification. FedIPR can also allow each client to embed their own backdoor-based watermark for contribution demonstration, but it may include harmful triggers and lack sanitization to prevent them. To ensure a fair comparison, we solely use backdoor-based watermarks in FedIPR. Besides, we also consider transforming sanitization mechanisms to be applied on the FL-trained model post-FL, without additional mechanisms employed during FL with watermark embedding. This represents an intuitive sanitizing approach that prioritizes efficiency, an efficiency-focused baseline (Efficiency-FB) that incurs no extra overhead during FL. Table VII shows the comparison of Sanitizer with FedIPR and Efficiency-FB in a consistent setting. While FedIPR achieves a watermark verification success rate of 99.25%, this stays the same for both harmful and harmless environments. For Efficiency-FB, while it does not incur extra burden during FL, its success rate for watermark verification is only 21.14%, whereas that for backdoor attacks remains high at 98.06%. On the contrary, Sanitizer successfully balances watermark detection, attack resistance, and model performance in our context.
Table VII: Effectiveness comparison of Sanitizer with two additional methods in consistent settings on CIFAR10 with ResNet18. Besides Sanitizer, neither method is able to satisfy all the properties in our context.
Approach WDR $\uparrow$ ASR $\downarrow$ ACC $\uparrow$
FedIPR 99.25% 99.25% 86.85%
Efficiency-FB 21.14% 98.06% 83.02%
Sanitizer 96.21% 11.10% 87.28%
5.9 Parameter Studies